Benny can dust the house in 11 hours by himself but when Mary helps it take 7 hours how long would it take Mary to do this on her own.
Answer by Boreal(15235)

in one hour, B does 1/11 of the house and together they do 1/7 of it in an hour. Mary takes x hours and does 1/x per hour.
(1/11)+(1/x)=1/7, equation for 1 hour
multiply by 77x
7x+77=11x
4x=77
x=19.75 hours, her time doing it.ANSWER
In 7 hours, Benny does 7/11 and Mary does 7/19.75